# Group class operations and homological conditions

### v1

@@ Abstract.
Kropholler’s operation ${\scriptstyle{{\bf LH}}}$ and Talelli’s operation $\Phi$ can be often used to formally enlarge the class of available examples of groups that satisfy certain homological conditions. In this paper, we employ this enlargement technique regarding two specific homological conditions. We thereby demonstrate the abundance of groups that (a) have virtually Gorenstein group algebras, as defined by Beligiannis and Reiten, and (b) satisfy Moore’s conjecture on the relation between projectivity and relative projectivity, that was studied by Aljadeff and Meir.
@@ 0. Introduction
Examples of this kind of homological conditions can be found in the realm of Gorenstein homological algebra, the relative homological theory which is based upon the classes of Gorenstein projective, Gorenstein flat and Gorenstein injective modules introduced in [23] and [25]. This theory has been developed rapidly and found interesting applications in algebraic geometry, representation theory and cohomological group theory. Even though important advances have been made recently by Šaroch and Št’ovíček [35], there are still many properties of the classical homological algebra notions that are not universally available for their Gorenstein homological algebra counterparts. The technique of using the operations ${\scriptstyle{{\bf LH}}}$ and $\Phi$ was employed in [22], in order to exhibit a very big class of groups that is convenient to work with, as far as these properties are concerned, in the sense that these properties are satisfied by the Gorenstein modules over the corresponding group algebras. In the present work, we provide two more instances of such an enlargement of the class of examples of groups that satisfy homological restrictions. The first of these examples deals with the notion of virtually Gorenstein algebras. This notion was introduced for Artin algebras by Beligiannis and Reiten in [5], in order to study the left-right symmetry of Gorenstein regularity; cf. [4, Theorem 8.7]. It provides a common generalization of Gorenstein regular algebras and algebras of finite representation type. The implications of the virtually Gorenstein condition to the homotopy (or the stable) categories of Gorenstein projective and Gorenstein injective modules are analysed in [3]; see also [11] and [37]. This notion was extended to arbitrary rings in [18]; it is proved therein that all weakly Gorenstein regular rings are virtually Gorenstein. Having fixed the commutative ring $k$, we may consider the class $\mathfrak{W}$ that consists of those groups $G$ for which $kG$ is weakly Gorenstein regular. If the class $\mathfrak{W}$ is non-empty, i.e. if $k$ is itself weakly Gorenstein regular, then $\mathfrak{W}$ contains all finite groups. For any $\mathfrak{W}$-group $G$, [18, Theorem A] implies that $kG$ is virtually Gorenstein. We now enlarge $\mathfrak{W}$ and let $\overline{\mathfrak{W}}$ be the smallest group class which contains $\mathfrak{W}$ and is both ${\scriptstyle{{\bf LH}}}$-closed and $\Phi_{inj}$-closed. Then, $\overline{\mathfrak{W}}$-groups admit a hierarchical description à la Kropholler, starting with $\mathfrak{W}$ and using (transfinite) induction based on the group class operation $\mathfrak{C}\mapsto{\scriptstyle{{\bf LH}}}\mathfrak{C}\cup\Phi_{inj}\mathfrak{C}$; this process is analysed in §$\S$1.IV. The enlargement of the group class of available examples that is advocated in this paper, is illustrated for the virtually Gorenstein property by the following result.
@@
This result is proved in §$\S 4$, using the description of the $\mbox{Ext}^{1}$-orthogonal classes of Gorenstein projective or Gorenstein injective modules over group algebras of groups that are obtained by performing Kropholler’s operation ${\scriptstyle{{\bf LH}}}$ (see also [29]) or Talelli’s operation $\Phi$ to certain group classes. The relation between these operations and $\mbox{Ext}^{1}$-orthogonality is analysed in §$\S 3$.
Another example of a homological condition that is well-behaved with respect to the operations ${\scriptstyle{{\bf LH}}}$ and $\Phi$ is provided by the relation between projectivity and relative projectivity for modules over group algebras. As an attempt to generalize the crucial step in the proof of Serre’s theorem on torsion-free groups of finite virtual cohomological dimension [10, Chapter VIII, §$\S$3], Moore has conjectured that a module over the group algebra of a group $G$ must be necessarily projective, provided that its restriction to a subgroup $H\subseteq G$ of finite index (that satisfies an additional necessary group-theoretic condition) is projective. Chouinard’s work [14] implies that all finite groups satisfy Moore’s conjecture. More generally, Aljadeff and Meir [1] have proved that Moore’s conjecture is satisfied by all ${\scriptstyle{{\bf LH}}}\mathfrak{F}$-groups. Building upon their work, in the present paper we improve the latter result as follows. Let $\overline{\mathfrak{F}}$ be the smallest class of groups which contains the class $\mathfrak{F}$ of finite groups and is ${\scriptstyle{{\bf LH}}}$-closed, $\Phi$-closed and $\Phi_{flat}$-closed. As before, $\overline{\mathfrak{F}}$-groups admit a hierarchical description starting with the class $\mathfrak{F}$ of finite groups and using (transfinite) induction based on the group class operation $\mathfrak{C}\mapsto{\scriptstyle{{\bf LH}}}\mathfrak{C}\cup\Phi\mathfrak{C}\cup\Phi_{flat}\mathfrak{C}$. The following result is proved in §$\S 5$.
@@
Here is a brief description of the contents of the paper: Following the preliminary Section 1, in Section 2, we fix the convenient universe of groups within which all Gorenstein homological algebra notions behave smoothly. In Section 3, we build on Kendall’s work [29] and describe the orthogonal modules with respect to the $\mbox{Ext}^{1}$-pairing of the Gorenstein classes over the group algebras of groups that are obtained by performing Kropholler’s or Talelli’s operation to group classes within the universe of groups described in Section 2. As an application of this description, we prove Theorem A in Section 4. Finally, in Section 5 we discuss Moore’s conjecture and combine the results by Aljadeff and Meir [1] with certain simple observations regarding the relation between projectivity and relative projectivity to prove Theorem B.
@@ 1. Preliminaries
I. Restriction and (co-)induction. Let $k$ be a commutative ring, $G$ a group and $H\subseteq G$ a subgroup. Restricting the action of $G$ on a $kG$-module $M$ to the subgroup $H$, we obtain the $kH$-module $\mbox{res}_{H}^{G}M$. In this way, we define the restriction functor $\mbox{res}_{H}^{G}$ from the category of $kG$-modules to that of $kH$-modules. This functor has both a left and a right adjoint. Its left adjoint is the induction functor $\mbox{ind}_{H}^{G}$; for any $kH$-module $N$ the induced $kG$-module $\mbox{ind}_{H}^{G}N$ is equal to $kG\otimes_{kH}N$, with $G$ acting on the left of $kG$. The right adjoint of restriction is the coinduction functor $\mbox{coind}_{H}^{G}$. Here, for any $kH$-module $N$ the coinduced $kG$-module $\mbox{coind}_{H}^{G}N$ is equal to $\mbox{Hom}_{kH}(kG,N)$, with $G$ acting on the right of $kG$. Any $kH$-module $N$ is contained as a direct summand in both $kH$-modules $\mbox{res}_{H}^{G}\mbox{ind}_{H}^{G}N$ and $\mbox{res}_{H}^{G}\mbox{coind}_{H}^{G}N$. For more details, the reader is referred to [10, Chapter III, §$\S$3]. Deriving the two adjunction isomorphisms, we obtain the Echmann-Shapiro isomorphisms: If $M$ is a $kG$-module and $N$ a $kH$-module, there are natural identifications
@@
for all $n\geq 0$. We also note that, in the special case where $H$ has finite index in $G$, the functors $\mbox{ind}_{H}^{G}$ and $\mbox{coind}_{H}^{G}$ are naturally isomorphic; cf. [10, Chapter III, Proposition 5.9].
II. Gorenstein modules. An acyclic complex of projective modules over a ring $R$ is called totally acyclic if it remains acyclic after applying the functor $\mbox{Hom}_{R}(\_\!\_,P)$ for any projective module $P$. A module $M$ is Gorenstein projective if it is a cokernel of a totally acyclic complex of projective modules; we denote by ${\tt GProj}(R)$ the class of these modules. Dually, an acyclic complex of injective modules is called totally acyclic if it remains acyclic after applying the functor $\mbox{Hom}_{R}(I,\_\!\_)$ for any injective module $I$. A module $N$ is Gorenstein injective if it is a kernel of a totally acyclic complex of injective modules; we denote by ${\tt GInj}(R)$ the class of these modules. An acyclic complex of flat modules is called totally acyclic if it remains acyclic after applying the functor $J\otimes_{R}\_\!\_$ for any injective right module $J$. A module $L$ is Gorenstein flat if it is a cokernel of a totally acyclic complex of flat modules; we denote by ${\tt GFlat}(R)$ the class of these modules. The Gorenstein modules defined above were introduced by Enochs, Jenda and Torecillas in [23] and [25]; see also Holm’s paper [27]. A distinguished class of Gorenstein flat modules has been introduced by Šaroch and Št’ovíček in [35]: A module $K$ is projectively coresolved Gorenstein flat if it is a cokernel of an acyclic complex of projective modules, which remains acyclic after applying the functor $J\otimes_{R}\_\!\_$ for any injective right module $J$; the class of these modules is denoted by ${\tt PGF}(R)$. As shown in [35, Theorem 4.4], projectively coresolved Gorenstein flat modules are Gorenstein projective, i.e. ${\tt PGF}(R)\subseteq{\tt GProj}(R)$.
@@
Non-trivial examples of cotorsion pairs are provided by Gorenstein modules. As shown in [35], there are complete hereditary cotorsion pairs $\left({\tt PGF}(R),{\tt PGF}(R)^{\perp}\right)$, $\left({\tt GFlat}(R),{\tt GFlat}(R)^{\perp}\right)$ and $\left({}^{\perp}{\tt GInj}(R),{\tt GInj}(R)\right)$ for any ring $R$. The situation regarding Gorenstein projective modules is more subtle: It is shown in [15] that $\left({\tt GProj}(R),{\tt GProj}(R)^{\perp}\right)$ is a hereditary cotorsion pair as well, but its completeness appears to depend upon set-theoretic assumptions.
@@
A group class $\mathfrak{C}$ is said to be closed under an operation $\mathcal{O}$ if $\mathcal{O}\mathfrak{C}=\mathfrak{C}$. The $\mathcal{O}$-closure of a group class $\mathfrak{C}$ is the smallest $\mathcal{O}$-closed group class containing $\mathfrak{C}$; it is the class consisting of those groups $G$ which are contained in $\mathfrak{D}$ for any $\mathcal{O}$-closed class $\mathfrak{D}$ that contains $\mathfrak{C}$. If the operation $\mathcal{O}$ is continuous, then there is an alternative hierarchical (bottom-to-top) description of the $\mathcal{O}$-closure of a group class $\mathfrak{C}$, that we shall now explicit. For any group class $\mathfrak{C}$, we define the classes $\mathfrak{C}_{\alpha}$ for any ordinal number $\alpha$, using transfinite induction, as follows:
@@
$0\longrightarrow M_{r}\longrightarrow\cdots\longrightarrow M_{1}\longrightarrow M_{0}\longrightarrow M\longrightarrow 0$
@@
$0\longrightarrow M\longrightarrow M^{0}\longrightarrow M^{1}\longrightarrow\cdots\longrightarrow M^{r}\longrightarrow 0,$
where $r$ is the dimension of $X$ and the $M_{i}$’s (resp. the $M^{i}$’s) are direct sums (resp. direct products) of $kG$-modules of the form $\mbox{ind}_{H}^{G}\mbox{res}_{H}^{G}M$ (resp. $\mbox{coind}_{H}^{G}\mbox{res}_{H}^{G}M$), for some ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{C}$-subgroup $H$ of $G$ with $\beta<\alpha$. We say that a group belongs to ${\scriptstyle{{\bf H}}}\mathfrak{C}$ if it belongs to ${\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{C}$ for some $\alpha$. Then, the class ${\scriptstyle{{\bf LH}}}\mathfrak{C}$ consists of those groups $G$, all of whose finitely generated subgroups $H$ are contained in some ${\scriptstyle{{\bf H}}}\mathfrak{C}$-subgroup $K\subseteq G$. The continuity of the operation ${\scriptstyle{{\bf LH}}}$ is proved in [22, Lemma 1.3]. If the class $\mathfrak{C}$ is subgroup-closed, then the classes ${\scriptstyle{{\bf H}}}\mathfrak{C}$ and ${\scriptstyle{{\bf LH}}}\mathfrak{C}$ are subgroup-closed as well; in this case, ${\scriptstyle{{\bf LH}}}\mathfrak{C}$ consists of those groups all of whose finitely generated subgroups are ${\scriptstyle{{\bf H}}}\mathfrak{C}$-groups.
@@
(ii) $\mbox{pd}_{kH}\mbox{res}_{H}^{G}M<\infty$ for any $\mathfrak{C}$-subgroup $H\subseteq G$.
See also [9, Definition 5.6]. The continuity of $\Phi$ is proved in [22, Lemma 1.4]. We shall also consider the version of the operation $\Phi$ for the flat (resp. injective) dimension of modules and define for any group class $\mathfrak{C}$ the class $\Phi_{flat}\mathfrak{C}$ (resp. $\Phi_{inj}\mathfrak{C}$) accordingly.

### v2

@@ Abstract.
Kropholler’s operation ${\scriptstyle{{\bf LH}}}$ and Talelli’s operation $\Phi$ can be often used to formally enlarge the class of available examples of groups that satisfy certain homological conditions. In this paper, we employ this enlargement technique regarding two specific homological conditions. We thereby demonstrate the abundance of groups that (a) have virtually Gorenstein group algebras, as defined by Beligiannis and Reiten, and (b) satisfy Moore’s conjecture on the relation between projectivity and relative projectivity, that was studied by Chouinard, Aljadeff, Cornick, Ginosar, Kropholler and Meir.
@@ 0. Introduction
Examples of this kind of homological conditions can be found in the realm of Gorenstein homological algebra, the relative homological theory which is based upon the classes of Gorenstein projective, Gorenstein flat and Gorenstein injective modules introduced in [24] and [26]. This theory has been developed rapidly and found interesting applications in algebraic geometry, representation theory and cohomological group theory. Even though important advances have been made recently by Šaroch and Št’ovíček [37], there are still many properties of the classical homological algebra notions that are not universally available for their Gorenstein homological algebra counterparts. The technique of using the operations ${\scriptstyle{{\bf LH}}}$ and $\Phi$ was employed in [23], in order to exhibit a very big class of groups that is convenient to work with, as far as these properties are concerned, in the sense that these properties are satisfied by the Gorenstein modules over the corresponding group algebras.
In the present work, we provide two more instances of such an enlargement of the class of examples of groups that satisfy certain (more or less restrictive) homological conditions. The first of these examples deals with the notion of virtually Gorenstein algebras. This notion was introduced for Artin algebras by Beligiannis and Reiten in [6], in order to study the left-right symmetry of Gorenstein regularity; cf. [5, Theorem 8.7]. It provides a common generalization of Gorenstein regular algebras and algebras of finite representation type. The implications of the virtually Gorenstein condition to the homotopy (or the stable) categories of Gorenstein projective and Gorenstein injective modules are analysed in [4]; see also [12] and [39]. The virtually Gorenstein property was studied for arbitrary rings in [19]; it is proved therein that all weakly Gorenstein regular rings are virtually Gorenstein. Having fixed the commutative ring $k$, we may consider the class $\mathfrak{W}$ that consists of those groups $G$ for which $kG$ is weakly Gorenstein regular. If the class $\mathfrak{W}$ is non-empty, i.e. if $k$ is itself weakly Gorenstein regular, then $\mathfrak{W}$ contains all finite groups. For any $\mathfrak{W}$-group $G$, [19, Theorem A] implies that $kG$ is virtually Gorenstein. We now enlarge $\mathfrak{W}$ and let $\overline{\mathfrak{W}}$ be the smallest group class which contains $\mathfrak{W}$ and is both ${\scriptstyle{{\bf LH}}}$-closed and $\Phi_{inj}$-closed. Then, $\overline{\mathfrak{W}}$-groups admit a hierarchical description à la Kropholler, starting with $\mathfrak{W}$ and using (transfinite) induction based on the group class operation $\mathfrak{C}\mapsto{\scriptstyle{{\bf LH}}}\mathfrak{C}\cup\Phi_{inj}\mathfrak{C}$; this process is analysed in §$\S$1.IV. The enlargement of the class of available examples, the central theme of this paper, is illustrated for the virtually Gorenstein property by the following result.
@@
This result is proved in §$\S 4$, using the description of the $\mbox{Ext}^{1}$-orthogonal classes of Gorenstein modules over group algebras of groups that are obtained by performing Kropholler’s operation ${\scriptstyle{{\bf LH}}}$ (see also [31]) or Talelli’s operation $\Phi$ to appropriate group classes. The relation between these operations and $\mbox{Ext}^{1}$-orthogonality is analysed in §$\S 3$.
Another example of a homological condition that is well-behaved with respect to the operations ${\scriptstyle{{\bf LH}}}$ and $\Phi$ is provided by the relation between projectivity and relative projectivity for modules over group algebras. As an attempt to generalize the crucial step in the proof of Serre’s theorem on torsion-free groups of finite virtual cohomological dimension [11, Chapter VIII, §$\S$3], Moore has conjectured that a module over the group algebra of a group $G$ must be necessarily projective, provided that its restriction to a subgroup $H\subseteq G$ of finite index (that satisfies an additional necessary group-theoretic condition) is projective. It follows from Chouinard’s work [15] that finite groups satisfy Moore’s conjecture. This result was extended to ${\scriptstyle{{\bf H}}}_{1}\mathfrak{F}$-groups in [1]. More generally, Aljadeff and Meir [2] have proved that Moore’s conjecture is satisfied by all ${\scriptstyle{{\bf LH}}}\mathfrak{F}$-groups. Building upon their work, in the present paper we improve the latter result as follows. Having fixed the commutative ring $k$, let $\overline{\mathfrak{F}}$ be the smallest class of groups which contains the class $\mathfrak{F}$ of finite groups and is ${\scriptstyle{{\bf LH}}}$-closed, $\Phi$-closed and $\Phi_{flat}$-closed. As before, $\overline{\mathfrak{F}}$-groups admit a hierarchical description starting with the class $\mathfrak{F}$ of finite groups and using (transfinite) induction based on the group class operation $\mathfrak{C}\mapsto{\scriptstyle{{\bf LH}}}\mathfrak{C}\cup\Phi\mathfrak{C}\cup\Phi_{flat}\mathfrak{C}$. The following result is proved in §$\S 5$.
@@
Here is a brief description of the contents of the paper: Following the preliminary Section 1, in Section 2, we fix the convenient universe of groups within which all Gorenstein homological algebra notions behave smoothly. In Section 3, we build on Kendall’s work [31] and describe the orthogonal modules with respect to the $\mbox{Ext}^{1}$-pairing of the Gorenstein classes over the group algebras of groups that are obtained by performing Kropholler’s or Talelli’s operation to group classes within the universe of groups described in Section 2. As an application of this description, we prove Theorem A in Section 4. Finally, in Section 5, we discuss Moore’s conjecture and combine the results by Aljadeff and Meir [2] with certain simple observations regarding the relation between projectivity and relative projectivity to prove Theorem B.
@@ 1. Preliminaries
I. Restriction and (co-)induction. Let $k$ be a commutative ring, $G$ a group and $H\subseteq G$ a subgroup. Restricting the action of $G$ on a $kG$-module $M$ to the subgroup $H$, we obtain the $kH$-module $\mbox{res}_{H}^{G}M$. In this way, we define the restriction functor $\mbox{res}_{H}^{G}$ from the category of $kG$-modules to that of $kH$-modules. This functor has both a left and a right adjoint. Its left adjoint is the induction functor $\mbox{ind}_{H}^{G}$; for any $kH$-module $N$ the induced $kG$-module $\mbox{ind}_{H}^{G}N$ is equal to $kG\otimes_{kH}N$, with $G$ acting on the left of $kG$. The right adjoint of restriction is the coinduction functor $\mbox{coind}_{H}^{G}$. Here, for any $kH$-module $N$ the coinduced $kG$-module $\mbox{coind}_{H}^{G}N$ is equal to $\mbox{Hom}_{kH}(kG,N)$, with $G$ acting on the right of $kG$. Any $kH$-module $N$ is contained as a direct summand in both $kH$-modules $\mbox{res}_{H}^{G}\mbox{ind}_{H}^{G}N$ and $\mbox{res}_{H}^{G}\mbox{coind}_{H}^{G}N$. For more details, the reader is referred to [11, Chapter III, §$\S$3]. Deriving the two adjunction isomorphisms, we obtain the Echmann-Shapiro isomorphisms: If $M$ is a $kG$-module and $N$ a $kH$-module, there are natural identifications
@@
for all $n\geq 0$. We also note that, in the special case where $H$ has finite index in $G$, the functors $\mbox{ind}_{H}^{G}$ and $\mbox{coind}_{H}^{G}$ are naturally isomorphic; cf. [11, Chapter III, Proposition 5.9].
II. Gorenstein modules. An acyclic complex of projective modules over a ring $R$ is called totally acyclic if it remains acyclic after applying the functor $\mbox{Hom}_{R}(\_\!\_,P)$ for any projective module $P$. A module $M$ is Gorenstein projective if it is a cokernel of a totally acyclic complex of projective modules; we denote by ${\tt GProj}(R)$ the class of these modules. Dually, an acyclic complex of injective modules is called totally acyclic if it remains acyclic after applying the functor $\mbox{Hom}_{R}(I,\_\!\_)$ for any injective module $I$. A module $N$ is Gorenstein injective if it is a kernel of a totally acyclic complex of injective modules; we denote by ${\tt GInj}(R)$ the class of these modules. An acyclic complex of flat modules is called totally acyclic if it remains acyclic after applying the functor $J\otimes_{R}\_\!\_$ for any injective right module $J$. A module $L$ is Gorenstein flat if it is a cokernel of a totally acyclic complex of flat modules; we denote by ${\tt GFlat}(R)$ the class of these modules. The Gorenstein modules defined above were introduced by Enochs, Jenda and Torecillas in [24] and [26]; see also Holm’s paper [29]. A distinguished class of Gorenstein flat modules has been introduced by Šaroch and Št’ovíček in [37]: A module $K$ is projectively coresolved Gorenstein flat if it is a cokernel of an acyclic complex of projective modules, which remains acyclic after applying the functor $J\otimes_{R}\_\!\_$ for any injective right module $J$; the class of these modules is denoted by ${\tt PGF}(R)$. As shown in [37, Theorem 4.4], projectively coresolved Gorenstein flat modules are Gorenstein projective, i.e. ${\tt PGF}(R)\subseteq{\tt GProj}(R)$.
@@
Non-trivial examples of cotorsion pairs are provided by Gorenstein modules. As shown in [37], there are complete hereditary cotorsion pairs $\left({\tt PGF}(R),{\tt PGF}(R)^{\perp}\right)$, $\left({\tt GFlat}(R),{\tt GFlat}(R)^{\perp}\right)$ and $\left({}^{\perp}{\tt GInj}(R),{\tt GInj}(R)\right)$ for any ring $R$. In fact, all three cotorsion pairs are cogenerated by suitable sets of modules. The situation regarding Gorenstein projective modules is more subtle: It is shown in [16] that $\left({\tt GProj}(R),{\tt GProj}(R)^{\perp}\right)$ is a hereditary cotorsion pair as well, but its completeness appears to depend upon set-theoretic assumptions.
@@
A group class $\mathfrak{C}$ is said to be closed under an operation $\mathcal{O}$ if $\mathcal{O}\mathfrak{C}=\mathfrak{C}$. The $\mathcal{O}$-closure of a group class $\mathfrak{C}$ is the smallest $\mathcal{O}$-closed group class containing $\mathfrak{C}$; it is the class consisting of those groups $G$ which are contained in $\mathfrak{D}$ for any $\mathcal{O}$-closed class $\mathfrak{D}$ that contains $\mathfrak{C}$. If the operation $\mathcal{O}$ is continuous, then there is an alternative hierarchical (bottom-to-top) description of the $\mathcal{O}$-closure of a group class $\mathfrak{C}$, that we now explicit. For any group class $\mathfrak{C}$, we define the classes $\mathfrak{C}_{\alpha}$ for any ordinal number $\alpha$, using transfinite induction, as follows:
@@
$0\longrightarrow M_{d}\longrightarrow\ldots\longrightarrow M_{1}\longrightarrow M_{0}\longrightarrow M\longrightarrow 0$
@@
$0\longrightarrow M\longrightarrow M^{0}\longrightarrow M^{1}\longrightarrow\ldots\longrightarrow M^{d}\longrightarrow 0,$
where $d$ is the dimension of $X$ and the $M_{i}$’s (resp. the $M^{i}$’s) are direct sums (resp. direct products) of $kG$-modules of the form $\mbox{ind}_{H}^{G}\mbox{res}_{H}^{G}M$ (resp. $\mbox{coind}_{H}^{G}\mbox{res}_{H}^{G}M$), for some ${\scriptstyle{{\bf H}}}_{\beta}\mathfrak{C}$-subgroup $H$ of $G$ with $\beta<\alpha$. We say that a group belongs to ${\scriptstyle{{\bf H}}}\mathfrak{C}$ if it belongs to ${\scriptstyle{{\bf H}}}_{\alpha}\mathfrak{C}$ for some $\alpha$. Then, the class ${\scriptstyle{{\bf LH}}}\mathfrak{C}$ consists of those groups $G$, all of whose finitely generated subgroups $H$ are contained in some ${\scriptstyle{{\bf H}}}\mathfrak{C}$-subgroup $K\subseteq G$. The continuity of the operation ${\scriptstyle{{\bf LH}}}$ is proved in [23, Lemma 1.3]. If the class $\mathfrak{C}$ is subgroup-closed, then the classes ${\scriptstyle{{\bf H}}}\mathfrak{C}$ and ${\scriptstyle{{\bf LH}}}\mathfrak{C}$ are subgroup-closed as well; in this case, ${\scriptstyle{{\bf LH}}}\mathfrak{C}$ consists of those groups all of whose finitely generated subgroups are ${\scriptstyle{{\bf H}}}\mathfrak{C}$-groups.
@@
(ii) there is an integer $n$, such that $\mbox{pd}_{kH}\mbox{res}_{H}^{G}M\leq n$ for any $\mathfrak{C}$-subgroup $H\subseteq G$.
See also [10, Definition 5.6]. The continuity of $\Phi$ is proved in [23, Lemma 1.4]. We shall also consider the version of the operation $\Phi$ for the flat (resp. injective) dimension of modules and define for any group class $\mathfrak{C}$ the class $\Phi_{flat}\mathfrak{C}$ (resp. $\Phi_{inj}\mathfrak{C}$) accordingly.